7. Consider the balanced equation : CH 4 (g)+ H 2 O(l) iff CO(g)+3H 2 (g) Initially, the concentration of the CH 4 is 0.200 M. At equilibrium, the concentration of the CO is 0.0500 M. Calculate K eq for this reaction.


1
Expert's answer
2021-02-28T06:15:54-0500

[CH4] = 0.2 - 0.05 = 0.15 M

[H2] = 3[CO] = 3"\times{}"0.05 = 0.15 M

K = "\frac{[CO\times[H_2]^3]}{[CH_4]} = \frac{0.05\times0.15^3}{0.15} = 1.125\times10^{-3} (\frac{mol^3}{l^3})"

Answer: K = 1.125"\times"10-3 "\frac{mol^3}{l^3}".